OpenStudy (anonymous):

the perimeter of a triangle is 47 centimeters. The longest side is 7 centimeters less than the sum of the other two sides. Twice the shortest side is 4 centimeters less than the longest side. Find the lengths of all sides.

OpenStudy (anonymous):

let the longest side be x,shortest side be z and the other side be y x+y+z=47 x=(y+z)-7 x=2z+4 now solve for x y and z

OpenStudy (anonymous):

a+b+c = 47cm let a be the longest side now a=b+c-7 shortest side be c 2c = a-4 c=(a-4)/2 a=b+c-7 a=b+(a-4)/2 (2a-a+4)/2 =b (a+4)/2=b a+b+(a-4)/2 = 47 cm (2a+2b+a-4) = 94cm (3a+2b-4)= 94cm solve further i think i did wrong some where have some work tell u later